OpenStudy (anonymous):

(square root of 6x-the square root of five) squared

OpenStudy (anonymous):

\[(\sqrt{6x}-\sqrt{5})^{2}\]

OpenStudy (anonymous):

\[(\sqrt(6x)-\sqrt(5) )^2\] ??

OpenStudy (anonymous):

thats what it looks like, i foiled and got the wrong answer

OpenStudy (anonymous):

apply the identity

OpenStudy (anonymous):

(a-b)^2